Grade 12Physical Chemistry

into bcc structure what fraction of edge is not covered by atoms

Aditya Singh Rathore
9 Years agoGrade 12

4 Answers

Vikas TU
9 Years ago
In BCC lattice of the atoms arrangement,
8 atoms are at corner and one at the body centre of te cubic crystal lattice.
And at edges only the portion which would not be covered would be due to the octahederal voids and that is => 
there are 12 octahederal voids and 8 tetrahederal voids and has a void fraction equals to => 32%.
And hence for for 20 voids it is 32% then for 12 voids it would be => 32*12/20 => 32*3/5 =69/5 => 13.8%
Karan
7 Years ago
Edge uncovered by atom = a-2r . Now we know that in bcc 4r = root 3 a . Therfore edge not covered = a ( 2-root3)/ 2
 Hence fraction of edge not covered = (2-root3) /2 
= (2-1.732)/2 = 0.268/2 = 0.134 
 Hope you understand !
Manoj
7 Years ago
2r/a......and put √3 a=4r......answer as packing is 0.866 and  unpacked 0.134....hope you enjoy this tremedoust answer
Rishi Sharma
5 Years ago
Dear Student,
Please find below the solution to your problem.

In BCC lattice of the atoms arrangement,
8 atoms are at corner and one at the body centre of te cubic crystal lattice.
And at edges only the portion which would not be covered would be due to the octahederal voids and that is =>
there are 12 octahederal voids and 8 tetrahederal voids and has a void fraction equals to => 32%.
And hence for for 20 voids it is 32% then for 12 voids it would be => 32*12/20 => 32*3/5 =69/5 => 13.8%

Thanks and Regards